OpenStudy (lukecrayonz):

Integral of (5/4)t(t^2+1)^3

OpenStudy (lukecrayonz):

\[\frac{ 5 }{ 4 }t(t^{2}+1)^{3}\]

OpenStudy (anonymous):

Try expanding the polynomial and then taking the antiderivative

OpenStudy (lukecrayonz):

@hartnn :(

OpenStudy (lukecrayonz):

Is it:\[\frac{ 5 }{ 8 }*\frac{ (t^{2}+1)^{4} }{ 4 }+C\]

hartnn (hartnn):

yes, thats correct